# Identification of Dynamic Nonlinear Panel Models

### v2

@@ 2.1 Model
In this section, we focus on the identification of dynamic binary choice models with endogenous covariates. Specifically, consider the following binary choice model:
@@
$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$ (5)
@@
where $Y_{it}\in\{0,1\}$ denotes the binary dependent variable for individual $i\in\{1,2,...\}$ and time $t\leq T$, $Z_{it}\in\mathcal{R}^{d_{z}}$ denotes exogenous covariates, $X_{it}\in\mathcal{R}^{d_{x}}$ denotes potentially endogenous covariates, $\alpha_{i}\in\mathcal{R}$ denotes the unobserved fixed effect for individual $i$, and $\epsilon_{it}$ denotes the unobserved time-varying error term for individual $i$ at time $t$. The objective is to identify the parameter $\theta_{0}:=(\beta_{0}^{\prime},\gamma_{0}^{\prime})^{\prime}$ using a short panel of observed variables $(Z_{it},X_{it},Y_{it})_{it}$, where the number of time periods $T\geq 2$ is fixed and finite.
The identification of model (5) has been explored in the literature under various assumptions. [^5] See Section 1 for more detailed discussions on related literature. Specifically, Manski (1987) exploits the following conditional stationarity (also referred to as “group homogeneity”) of $\epsilon_{it}$ to achieve identification:
$\epsilon_{is}\mid Z_{is},Z_{it},X_{is},X_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{is},Z_{it},X_{is},X_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$ (6)
The above condition does not impose parametric restrictions on the distributions of $\epsilon_{it}$ and allows for dependence of $\epsilon_{it}$ over time. However, condition (2) does impose substantial restriction on the dependence between $(Z_{it},X_{it})$ and the time-varying error $\epsilon_{it}$. For instance, suppose $\mathbb{E}[\left.\epsilon_{it}\right|Z_{is},Z_{it},X_{is},X_{it}]=Z_{it}^{\prime}\eta_{1}+X_{it}^{\prime}\eta_{2}$, then condition (2) fails in general. Hence, condition (2) can also be interpreted as a form of exogeneity condition on $\epsilon_{it}$ and all covariates $(Z_{it},X_{it})$. Importantly, in a dynamic setting where $X_{it}$ includes the lagged dependent variable $Y_{i,t-1}$, then condition (2) is no longer a justifiable assumption to work with.
Our paper aims to study identification of $\theta_{0}$ robust to endogeneity of $X_{it}$, where $X_{it}$ can be arbitrarily dependent with $(\alpha_{i},\epsilon_{it})$. Before introducing our assumptions, we first provide several applications to illustrate this endogeneity.
Example 1 (Dynamic Models).
Khan, Ponomareva, and Tamer (2023) studies the following dynamic model with one lagged dependent variable:
$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+Y_{it-1}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$
under the stationarity assumption only given exogenous covariates $Z_{it}$: $\epsilon_{is}\mid Z_{is},Z_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}$. In this model, the full stationarity condition may naturally fail due to its dynamic nature. For example, suppose $\epsilon_{it}$ is i.i.d. across time and independent of ($(\{Z_{it}\}_{t=1}^{T},\alpha_{i})$. Given $Z_{is}=Z_{it}=z,\alpha_{i}=a$, it is clear that the conditional distribution of $\epsilon_{it}\mid Y_{it-1},Y_{it},z,a$ will be different from $\epsilon_{it+1}\mid Y_{it-1},Y_{it},z,a\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it+1}$, since $Y_{it}$ is informative for $\epsilon_{it}$.
Example 2 (Omitted Variable).
Consider the true model is given as $Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+S_{it}\eta_{0}+\alpha_{i}+v_{it}\geq 0\}$, where $v_{it}$ is independent of $(\alpha_{i},\{Z_{it},X_{it},S_{it}\}_{t=1}^{T})$. However, the variable $S_{it}$ is omitted and we instead estimate the following model:
$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$
where $\epsilon_{it}=v_{it}+S_{it}\eta_{0}$. The covariate $X_{it}$ can be correlated with $\epsilon_{it}$ through the omitted variable $S_{it}$, provided that the omitted variable $S_{it}$ is correlated with $X_{it}$. Then the conditional distribution of $\epsilon_{it}\mid(Z_{is},Z_{it},X_{is},X_{it},\alpha_{i})$ may differ from $\epsilon_{is}\mid(Z_{is},Z_{it},X_{is},X_{it},\alpha_{i})$ since the distribution of $S_{it}\mid X_{it}$ could be different from the distribution $S_{is}\mid X_{is}$ when $X_{it}\neq X_{is}$.
Example 3 (Measurement Error).
Consider the true model is given as $Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{*^{\prime}}\gamma_{0}+\alpha_{i}+v_{it}\geq 0\}$, where $v_{it}$ is independent of $\left(\alpha_{i},\{Z_{it},X_{it}^{*}\}_{t=1}^{T}\right)$. However, the covariate $X_{it}^{*}$ may not be observed and we only observe its measurement $X_{it}$ which may be subject to a measurement error $X_{it}=X_{it}^{*}+e_{it}$. Suppose we estimate the following model using the measurement $X_{it}$ as a regressor:
$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$
where $\epsilon_{it}=v_{it}-e_{it}^{\prime}\gamma_{0}$. The error term $\epsilon_{it}$ would be correlated with the measurement $X_{it}$ through the measurement error $e_{it}$. Similarly, the conditional stationarity in (2) could fail since $X_{it}$ is informative for $e_{it}$ such that the conditional distribution of $\epsilon_{it}\mid z,X_{is},X_{it},\alpha_{i}$ would vary over time given different covariates $X_{is}\neq X_{it}$ while holding $Z_{is}=Z_{it}=z$.
Motivated by the above examples, this paper aims to establish identification of $\theta_{0}$, while being robust to endogeneity of $X_{it}$ with respect to unobserved factors $(\alpha_{i},\epsilon_{it})$. Below we propose an approach that accommodates various types of endogeneity, regardless of whether the endogenous covariates $X_{it}$ are discrete lagged outcome variables (as in Example 1) or continuously-valued covariates with contemporary endogeneity as in other examples. To achieve this, we introduce the following partial conditional stationarity assumption:
@@
The conditional distribution of $\epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}$ is stationary over time:
$\epsilon_{is}\mid Z_{is},Z_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$
Assumption 1 only assumes the stationarity given the covariate $(Z_{is},Z_{it})$, but does not impose any restrictions on the endogenous covariate $X_{it}$. The standard full stationarity condition (2) is nested in Assumption 1, and it is in general stronger than Assumption 1. The main distinction is that Assumption 1 allows $X_{it}$ to be arbitrarily dependent with both the fixed effect $\alpha_{i}$ and the time-changing error $\epsilon_{it}$. In addition, their correlation can be different across time. Assumption 1 also shares similar features with condition (2), in that it does not impose any distributional assumption on $\epsilon_{it}$ and allows $\epsilon_{it}$ to be correlated across time. The standard i.i.d. assumption of $\epsilon_{it}$ across time or independence of $\epsilon_{it}$ with $(\alpha_{i},Z_{it},X_{it})$ are also nested in Assumption 1 as special cases.
2.2 Identification Strategy
We now explain our key identification strategy based on partial stationarity. While it applies more generally, we first illustrate the key idea in the specific context of the dynamic binary choice model. Sections 3-6 provide generalizations of the idea to various models.
Let $v_{it}:=-(\epsilon_{it}+\alpha_{i})$, $W_{it}:=(Z_{it},X_{it})$, and $W_{ist}:=(W_{is},W_{it})$. The conditional choice probability is given as follows:
$\Pr\left(Y_{it}=1\mid W_{ist}=w_{st}\right)=\Pr\left(v_{it}\leq w_{t}^{\prime}\theta_{0}\mid W_{ist}=w_{st}\right).$
When assuming all covariates are exogenous under condition (2), variation in the choice probability is solely due to variation in the covariate index $w_{t}^{\prime}\theta_{0}$. [^6] Condition 2 implies the stationarity of $v_{it}$: $v_{is}\mid W_{ist}\stackrel{{\scriptstyle d}}{{\sim}}v_{it}\mid W_{ist}$. Therefore, as shown in Manski (1987), the identified set for $\theta_{0}$ is characterized by the condition: an increase in the value of $w_{t}^{\prime}\theta_{0}$ over time implies an increase in the choice probability over time.
However, Manski (1987)’s method is not applicable to our setting with Assumption 1, given that the endogenous regressor $X_{it}$ can be correlated with $v_{it}$. When the covariate $X_{it}$ changes, the conditional distribution of $v_{it}$ might potentially change as well. Variation in the choice probability contains a mixture of changes in the covariate index and the conditional distribution of $v_{it}$. Hence, it is not feasible to derive identifying restrictions on $\theta_{0}$ merely by examining the intertemporal variation in the choice probability.
Our analysis relies on the partial stationarity condition in Assumption 1, which implies the stationarity of $v_{it}$ given $Z_{ist}:=(Z_{is},Z_{it})$:
$v_{is}\mid Z_{ist}\stackrel{{\scriptstyle d}}{{\sim}}v_{it}\mid Z_{ist}.$
The identification strategy proceeds in two steps. We first derive lower and upper bounds for the distribution of $v_{it}\mid Z_{ist}$ at each period $t$. The derived bounds at each period $t$ are functions of observed variables and the parameter $\theta_{0}$. Since the distribution of $v_{it}\mid Z_{ist}$ is the same over time, then the bounds of the distribution across time should have intersections. This restriction serves as the identifying restriction for $\theta_{0}$. Next, we show how to derive bounds on the distribution of $v_{it}\mid Z_{ist}$.
Let $F_{v_{t}\mid W_{st}}(\cdot\mid w_{st})$ denote the conditional CDF of $v_{it}$ given $W_{ist}=w_{st}$. For any fixed point $a\in\mathcal{R}$, we first derive bounds for the conditional distribution $F_{v_{t}\mid W_{st}}(a\mid w_{st})$, and then establish bounds on $F_{v_{t}\mid Z_{st}}\left(a\mid z_{st}\right)$ by taking expectation over the covariate $X_{ist}$ given $Z_{ist}=z_{st}$.
Lower bound. The conditional choice probability given $W_{ist}=w_{st}$ is given as
$\Pr\left(Y_{it}=1\mid W_{ist}=w_{st}\right)=\Pr(v_{it}\leq w_{t}^{\prime}\theta_{0}\mid W_{ist}=w_{st})=F_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta_{0}\mid w_{st}).$
For any fixed point $a$, by the monotonicity of a distribution, $F_{v_{t}\mid W_{st}}(a\mid w_{st})$ can be bounded below using the observed choice probability:
$F_{v_{t}\mid W_{st}}(a\mid w_{st})\geq\Pr(Y_{it}=1\mid W_{ist}=w_{st})\mathbbm{1}\{w_{t}^{\prime}\theta_{0}\leq a\}.$
The above inequality holds since either $a$ is larger than the covariate index $w_{t}^{\prime}\theta_{0}$ so we know $F_{v_{t}\mid W_{st}}(a\mid w_{st})\geq F_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta_{0}\mid w_{st})$ or the right hand side is equal to zero when $a$ is smaller than $w_{t}^{\prime}\theta_{0}$.
As Assumption 1 is on the distribution of $v_{it}$ given $Z_{ist}$, we can establish the lower bound for $F_{v_{t}\mid Z_{st}}(a\mid z_{st})$ by taking expectation over covariates $X_{ist}$ given $Z_{ist}=z_{st}$:
@@ 2.2 Identification Strategy
$=\Pr(Y_{it}=1,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\leq a\mid z_{st}).$
The above lower bound only depends on observed variables $(W_{ist},Y_{it})$ and $\theta_{0}$, so it is identified up to the parameter $\theta_{0}$. This bound will help construct identifying restrictions on $\theta_{0}$.
Upper bound. The idea of constructing the upper bound is similar. For any $a\in\mathcal{R}$, the conditional distribution $F_{v_{t}\mid W_{st}}$ can be bounded above:
$F_{v_{t}\mid W_{st}}(a\mid w_{st})\leq\Pr(Y_{it}=1\mid W_{ist}=w_{st})\mathbbm{1}\{w_{t}^{\prime}\theta_{0}\geq a\}+\mathbbm{1}\{w_{t}^{\prime}\theta_{0}<a\}.$
Then, taking expectation over covariates $X_{ist}$ given $Z_{ist}=z_{st}$ leads to the upper bound for $F_{v_{t}\mid Z_{st}}(a\mid z_{st})$:

### v3

@@ 2.1 Model
We start with an illustration of our general identification strategy in the binary choice setting. Specifically, consider the following binary choice model:
@@
$Y_{it}=\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\right\},$ (5)
@@
where a panel of data $\left(Y_{it},Z_{it},X_{it}\right)$ are observed across individuals $i=1,...,n$ and time periods $t=1,...,T$. Here $Y_{it}$ denotes the binary outcome variable, $Z_{it}$ and $X_{it}$ denote two types of observed covariates (the difference between $Z_{it}$ and $X_{it}$ will be clarified below), $\alpha_{i}$ denotes the unobserved fixed effect for individual $i$, and $\epsilon_{it}$ denotes the unobserved time-varying error term for individual $i$ at time $t$. The objective is to identify the unknown parameter $\theta_{0}:=\left(\beta_{0}^{{}^{\prime}},\gamma_{0}^{{}^{\prime}}\right)^{{}^{\prime}}$, with $\beta_{0}\in\mathcal{R}^{d_{z}}$ and $\gamma_{0}\in\mathcal{R}^{d_{x}}$. We focus on the short panel setting, where the number of individuals $n$ is considered to be larg while the number of time periods $T\geq 2$ is fixed and finite.
Write $Z_{i}:=\left(Z_{i1},...,Z_{iT}\right)$ and $X_{i}:=\left(X_{i1},...,X_{iT}\right)$. Throughout this paper, we will refer to $Z_{i}$ as “exogenous covariates”, and refer to $X_{i}$ as “endogenous covariates”. The exact difference between $Z_{i}$ and $X_{i}$, as well as the precise meaning of “exogeneity” versus “endogeneity” in our context, are formalized through the following “partial stationarity” assumption: the word “partial” emphasizes that this assumption is imposed on only a part of the observed covariates $Z_{i}$, but not on $X_{i}$.
@@
The conditional distribution of $\epsilon_{it}\mid Z_{i},\alpha_{i}$ is stationary over time, i.e.,
$\epsilon_{it}\mid Z_{i},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{is}\mid Z_{i},\alpha_{i}\quad\forall t,s=1,...,T.$
Assumption 1 essentially requires that the (conditional) distribution of $\epsilon_{it}$ stays the same across all time periods $t=1,...,T$ even if $Z_{i}$ realize to different values. To illustrate, suppose that there are only two periods $t=1,2$, and that $Z_{i1},Z_{i2}$ realize to two values $z_{1},z_{2}$, respectively, with $z_{1}<z_{2}$. Then Assumption 1 requires that $\epsilon_{i1}$ and $\epsilon_{i2}$ still have the same (conditional) distributions: in particular, $\epsilon_{i1}$ cannot be stochastically smaller (or larger) than $\epsilon_{i2}$ because of $z_{1}<z_{2}$. Hence, Assumption 1 can be thought as a definition of the “exogeneity” of the covariates $Z_{it}$ in our context.
In contrast, Assumption 1 imposes no such restrictions on the (potentially) endogenous covariates $X_{i}$. In fact, since $X_{i}$ does not appear in Assumption 1 at all, here we are completely agnostic about the dependence structure between $\epsilon_{i1},...,\epsilon_{iT}$ and $X_{i}$: in particular, the conditional distribution of $\epsilon_{it}$ is allowed to vary across $t$ arbitrarily for any particular realization of $X_{i}$. As a result, different forms of endogeneity in $X_{i}$ can be incorporated under our framework in a unified manner, as we illustrate in the examples below.
Example 1 (Dynamic Effects via Lagged Outcomes).
Consider the following “AR(1)-type” dynamic binary choice model studied in Khan, Ponomareva, and Tamer (2023, KPT thereafter):
$Y_{it}=\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+Y_{i,t-1}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\right\},$
which is a special case of our model with $X_{it}$ set to be the one-period lagged binary outcome variable $Y_{i,t-1}$. Here, $X_{it}$ is endogenous since $X_{it}\equiv Y_{i,t-1}$ and $\epsilon_{i,t-1}$ is by construction positively correlated with $Y_{i,t-1}$ for any $t$, and thus the distribution of $\epsilon_{it}$ cannot be stationary across $t$ when conditioned on the realizations of $Y_{i0},...,Y_{i,T-1}$. For example, given $Y_{i0}=Y_{i1}=1,$ $Y_{i2}=0$ (and $Z_{i},\alpha_{i}$), the conditional distribution of $\epsilon_{i1}$ will naturally be different from that of $\epsilon_{i2}$. To obtain identification under the endogeneity of $Y_{i,t-1}$, KPT imposes the stationarity of $\epsilon_{it}$ conditional on the exogenous covariates $Z_{i}$ only, which coincides with our “partial stationarity” condition (Assumption 1) when specialized to their setting.
A natural generalization of the AR(1) model above in KPT is the following “AR($p$)” model, which is again a special case of our model with $X_{it}$ taken to be the vector of $p$ lagged outcomes $Y_{i,t-1},...,Y_{i,t-p}$:
$Y_{it}=\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+\sum_{j=1}^{p}Y_{i,t-j}\gamma_{j}+\alpha_{i}+\epsilon_{it}\geq 0\right\}.$
Similarly, $X_{it}$ is endogenous here due to dependence on $\epsilon_{i,t-1},...,\epsilon_{t-p}$, which can again be handled in our framework under the “partial stationarity” assumption. While it is not clear how the identification results in KPT can be easily generalized to the AR($p$) model above, we show in the next subsection how our identification strategy provides a simple and unified approach to derive moment inequalities regardless of the exact specifications of $X_{it}$.
Example 2 (Contemporarily Endogenous Covariates).
Alternatively, consider the following binary choice model with contemporarily endogenous covariates:
$Y_{it}$ $=\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\right\},$
$X_{it}$ $=\phi\left(Z_{it},u_{it}\right)$
where $\phi$ is an unknown “first-stage” function and the “first-stage error” $u_{it}$ is allowed to be arbitrarily correlated with $\epsilon_{it}$. For example, $X_{it}$ may be a “price-type” variable that is strategically chosen by a decision maker after observing the current-period error $\epsilon_{it}$, which generates contemporary dependence between $X_{it}$ and $\epsilon_{it}$. Even though contemporary endogeneity of this type is very different in nature from the dynamic endogeneity discussed in the previous example, it also induces non-stationarity of $\epsilon_{it}$ when conditioned on $X_{i}$: for example, if $X_{it}$ and $\epsilon_{it}$ are positively correlated, then, conditional on $X_{i1}<X_{i2}$, it is unreasonable to assume the distribution of $\epsilon_{i1}$ is the same as $\epsilon_{i2}$. That said, such type of contemporary endogeneity can also be handled in our framework under the “partial stationarity” condition (Assumption 1).
Remark 1 (Combination of Dynamic and Contemporary Endogeneity).
We separately discussed two types of endogenous covariates, dynamic covariates (lagged outcome variables) and contemporarily endogenous covariates, in the two examples above, but our identification strategy also applies if both types of endogenous covariates are present together, since our identification strategy works generally under “partial stationarity”, which does not impose or exploit any restrictions on the form of endogeneity between $\epsilon_{it}$ and $X_{i}$.
Remark 2 (Full Stationarity as Special Case).
Obviously, the standard “full stationarity” condition (2) is nested under “partial stationarity” condition (Assumption 1) as a special case, where the endogenous covariate $X_{it}$ contains no variables. Hence, “full stationarity” is in general stronger than “partial stationarity”.
Remark 3 (Focus on Time-Varying Endogeneity).
Technically, our partial stationarity condition also allows some endogeneity between $\epsilon_{it}$ and $Z_{i}$, as long as such endogeneity is time-invariant. This is because Assumption 1 is stated conditional on the full vector $Z_{i}=\left(Z_{i1},...,Z_{iT}\right)$ and the time-invariant fixed effect $\alpha_{i}$. Hence, as long as the conditional distribution of $\epsilon_{it}$ depends on $Z_{i1},...,Z_{iT}$ and $\alpha_{i}$ in a time-invariant manner, the stationarity of $\epsilon_{it}$ can still hold. That said, since in empirical applications we are mostly interested in “time-varying endogeneity”, such as the dynamic and contemporary endogeneity discussed in the examples above, in this paper we refer to $Z_{i}$ as “exogenous” even though it may be endogenous in a time-invariant manner, and only call $X_{i}$, which features time-varying endogeneity, the “endogenous” covariates.
Remark 4 (Pairwise Version of Partial Stationarity).
In Assumption 1, we impose partial stationarity of $\epsilon_{it}$ conditional on $Z_{it}$ from all periods $t=1,...,T$. Alternatively, we could impose partial stationarity in a “pairwise” version, conditional on $\left(Z_{it},Z_{is}\right)$ from any pair of time periods $\left(t,s\right)$ only:
@@
Clearly, the “pairwise” version is equivalent to the “all-periods” version when $T=2$, but is weaker when $T\geq 3$. Our identification strategy applies under both versions of partial stationarity, though the identification results and the corresponding proofs have slightly different representations. Essentially, conditioning on all-period covariate realizations would be replaced with conditioning the realizations in any specific pair of period. See Remark 5 at the end of Section 2.2 for a follow-up discussion.
2.2 Key Identification Strategy
We now explain our key identification strategy based on partial stationarity.
Write $v_{it}:=-\left(\epsilon_{it}+\alpha_{i}\right)$ so that model (5) can be rewritten as
$Y_{it}=\mathbf{\mathbbm{1}}\left\{v_{it}\leq Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\right\}.$
For any constant $c\in\mathcal{R}$, consider first the event
$Y_{it}=1\text{ and }Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c.$
Whenever the event above happens, we must have $v_{it}\leq Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c$, implying that $v_{it}\leq c$. Formally, the above can be summarized by the following inequality:
$Y_{it}\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c\right\}$ $=\mathbf{\mathbbm{1}}\left\{v_{it}\leq Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\right\}\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c\right\}$ (7)
$\leq\mathbf{\mathbbm{1}}\left\{v_{it}\leq c\right\}.$
Symmetrically, we can also consider the “flipped” event
$Y_{it}=0\text{ and }Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\geq c,$
which implies $v_{it}>c$, and similarly
$\left(1-Y_{it}\right)\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\geq c\right\}=$ $\mathbf{\mathbbm{1}}\left\{v_{it}>Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\right\}\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\geq c\right\}$